OpenStudy (halfdeafdarling):

Hey anyone able to help with a square root problem? I am confused with how they simplified something and need help clearing it up.

OpenStudy (halfdeafdarling):

Okay so they say simplify \[\sqrt{-72}\] while going through i do the -1 * -72 and -1 = i etc. but the answers they give is like -6squared by two

OpenStudy (leozap1):

6i√2

OpenStudy (leozap1):

That's what I got at least. Did you get the same thing?

OpenStudy (halfdeafdarling):

\[so far I have \sqrt{9} * \sqrt{8} which then I simplify\]

OpenStudy (halfdeafdarling):

The thing is how did you get that? I am confused

OpenStudy (anonymous):

\[\sqrt{-72} = \sqrt{36*2*-1} = 6\sqrt{2}\sqrt{-1} = 6\sqrt{2}i\]

OpenStudy (halfdeafdarling):

oh okay --- that shows how that got that -- Its confusing the way they have it is you multiply just two numbers.

OpenStudy (leozap1):

The same way Zelda just said.

OpenStudy (anonymous):

@leozap1 Haha. :)

OpenStudy (anonymous):

I saw that lol. ;)

OpenStudy (leozap1):

Lol.

OpenStudy (halfdeafdarling):

thanks zelda -- how do you figure that you do that with the more expanded multiplication?

OpenStudy (anonymous):

You said 6. So I know 36 = 6^2, and I know 2*36 = 72.. so I just went backwards.

OpenStudy (halfdeafdarling):

and may I ask if i have another question right? It is \[\sqrt{-121}\] and i got 11i is that correct? -------- oooh and okay that makes more sense on that question ill have to keep that in mind for future reference

OpenStudy (anonymous):

Yea, 11i is right.

OpenStudy (halfdeafdarling):

okay thanks -- i figured that one out easy enough since its just 11 * 11 but they never shown me how to do the more expanded multiplication. Again thanks for the help zelda